A particle moves along the x-axis so that at time t≥0 its position is given by x(t) = -t^2 - 13t + 5. Determine the acceleration of the particle at t = 5.


1
Expert's answer
2021-02-24T06:54:59-0500

Consider the position function "x(t)=-t^2-13t+5"


Differentiate "x(t)" with respect to "t" in order to find the velocity function "v(t)" as,


"v(t)=\frac{d}{dt}[x(t)]"


"=\frac{d}{dt}(-t^2-13t+5)"


"=-(2t)-13(1)"


"=-2t-13"


Differentiate "v(t)" with respect to "t" in order to find the acceleration function "a(t)" as,


"a(t)=\frac{d}{dt}[v(t)]"


"=\frac{d}{dt}(-2t-13)"


"=-2(1)-0"


"=-2"


Here, "a(t)=-2" is a constant function, so the acceleration is constant for "t\geqslant0" .


Therefore, the acceleration of the particle at "t=5" is "a(5)=-2" .